OpenStudy (anonymous):

if cos theta=-3/5 and 90

zepdrix (zepdrix):



OpenStudy (anonymous):

not really but if you could explain that would be awesome! Or if you could show me the steps on how to work out sin2theta i should be able to work my way around it. i know how to work out sin theta but i just have no clue on how to work out sin 2 theta....

zepdrix (zepdrix):

You know how to work out the \(\sin \theta\)? Ok cool :) Just so we're clear, is it \(\sin 2\theta\) or \(\sin ^2\theta\)?

OpenStudy (anonymous):

it's sin2θ

zepdrix (zepdrix):

For \(\sin 2\theta\) we'll want to simply apply our Double Angle Formula for Sine.\[\large \sin 2\theta = 2\sin \theta \cdot \cos \theta\]

zepdrix (zepdrix):

Shouldn't be too bad from here, since we were GIVEN cosine, and you seem to have already figured out sine, so just plug them in! :)

OpenStudy (anonymous):

i have but i got the wrong answer, i thought i was using the wrong formula... :/

OpenStudy (anonymous):

ohh ive got it its okay my calc was on the worng settting

zepdrix (zepdrix):

\[\large \sin \theta = \frac{4}{5}, \qquad \cos \theta=-\frac{3}{5}\] \[\large 2\sin \theta \cdot \cos \theta \quad = \quad 2 \left(\frac{4}{5}\right)\left(-\frac{3}{5}\right)\]Oh hah :) that's silly

OpenStudy (anonymous):

hahaha just something i would do! :D but thank you !

OpenStudy (anonymous):

just another quire, if i had \[\cos2\theta=2/3\] how would i go about finding sintheta

zepdrix (zepdrix):

Hmmmmm